It takes 23 hours 56 minutes and 4 seconds for the earth to make one revolution 7.29x10-5 rad/s Your recepir nos is 155-2480 Previous Tries Assume the earth is spherical. Relative to someone on the rotation axis, what is the linear speed of an object on the surface if the redius vector from the center of the earth to the object makes an angle of 7.0 with the axis of rotation. The radius of the earth is 637.0)km. 4.39x102 m/s You are correct. Your receipt no. is 155-6429 Previous Tries eleration of the object on the surface of the earth in the previous problem? What is the acceleration of an object undergoing uniform circular motion?
Answer #1


acceleration a = r*w^2 = R*sin71*w^2

R = radius of earth

r = perpendicular distance from axis of rotation to object on earth at 71o

a = 6.37*10^6*sin71*(7.29*10^-5)^2


a = 0.032 m/s^2